the directions are solve each equation
the problem is n+3 2/3 = 5 3/8
this is how far i have gotten
n+ 3 2/3 = 5 3/8
-3 2/3 -3 2/3
n= 5 3/8- 3 2/3

Answer by stanbon(75887)

the problem is n+3 2/3 = 5 3/8
---
n + 11/3 = 43/8
---
Multiply thru by 24 to get:
24n + 8*11 = 3*43
---
24n = 3*43-88
24n = 41
n = 41/24
-----------------
Cheers,
Stan H.